OpenStudy (anonymous):

find the linear factors of a) 6z^2 - z^2 - 2z b) z^4 -6z^2 +5 find the linear factors of a) 6z^2 - z^2 - 2z b) z^4 -6z^2 +5 @Mathematics

OpenStudy (anonymous):

a) z(5z - 2)

OpenStudy (anonymous):

Working on the second answer....

OpenStudy (anonymous):

Do you think you should use the method for factorising a quadratic expression?

OpenStudy (anonymous):

yeahh !

OpenStudy (anonymous):

Oh yeah, got the answer - b) (z^2 - 5)(z^2 - 1)

OpenStudy (anonymous):

Hope my answer was useful. Want to know the workings?

OpenStudy (anonymous):

yahh i would like to know the workings! hahah! >D thank you so much !

OpenStudy (anonymous):

a) 6z^2 - z^2 - 2z = 5z^2 - 2z Take common factor(z). z(5z - 2) b) z^4 - 6z^2 + 5 Sum = -6 Product = 5 z^4 - z^2 - 5z^2 + 5 = z^2(z^2 - 1) - 5(z^2 - 1) = (z^2 - 5)(z^2 - 1)

OpenStudy (anonymous):

Oh, come on, no medals?? And I took so much effort in typing out that answer!!

OpenStudy (anonymous):

Ah, that's better.